# Strategic Informed Trading and the Value of Private Information

### v1

@@ 1. The Equilibrium
We now present the model and construct the equilibrium. To both isolate the effects of price impact internalization and asymmetric information, and to keep the presentation/notation as simple as possible, we consider a simplified model with only one risky tradeable asset whose payoff has zero mean and unit variance, and where the initial endowments are at Pareto optimality absent private information [^6] Even though traders have CARA preferences, equilibrium quantities under price-impact do depend on the traders’ initial endowments. Intuitively, this is because it is insider’s trade off her initial position that impacts equilibrium prices. Prices then depend on the insider’s initial endowment, and through market-clearing also on the uniformed traders’ initial position. Assuming initial endowments are Pareto optimal absent private information turns off the channel where welfare differences arise due to hedging demands based on the initial position, independently of private information, and allows us to focus solely on information and internalization effects. While the main discussion of the paper is developed under this assumption, all the equilibrium formulas are stated in the Appendices for general initial positions consistent with the outstanding supply.. The single asset model, and all the main results, are generalized in Appendix A removing the afore-mentioned assumptions. With the exception of Proposition 1.5, proofs of all statements made in this section are given for the general case in Appendix B.
@@
The model has one period. The risky asset has terminal payoff $X\sim N(0,1)$ and has positive supply $\Pi>0$. The risk-less asset is in 00 net supply, with price normalized to $1$. Following the literature (see Subrahmanyam (1991); Spiegel and Subrahmanyam (1992); Grossman (1976); Grossman and Stiglitz (1980) amongst many others), there is an insider $I$ who at time 00 obtains a private signal $G$ which is a noisy version of $X$, taking the form
(1) $G=X+Z_{I};\qquad Z_{I}=\frac{1}{\sqrt{p_{I}}}\mathcal{E}_{I},$
where $\mathcal{E}_{I}\sim N(0,1)$ is independent of $X$ and $p_{I}>0$ is the signal precision. There is also a mass of uniformed traders who do not receive a private signal, but who in the equilibrium established below, will receive a market signal through the time 00 price. In contrast to Nezafat and Schroder (2023), we assume the uninformed traders are price-takers, and following convention, we consider a representative agent $U$, hereafter called the uniformed trader. We assume both traders have exponential preferences with respective risk tolerances $\alpha_{I}$, $\alpha_{U}$. Lastly, there are liquidity providers (also called noise traders), denoted by $N$, with exogenous demand
@@
where $\mathcal{E}_{N}\sim N(0,1)$ is independent of both $X$ and $\mathcal{E}_{I}$. $p_{N}$ measures the noise trader demand precision, and, as mentioned in Kovalenkov and Vives (2014b) can be thought of as a measure of the volume of the price inelastic demand. Traders $I$ and $U$ are endowed with (constant) share positions $\left\{\pi_{i,0}\right\}$ which are Pareto optimal absent private information [^7] Since the uninformed agent, being a price taker, can be seen as a representative agent for a group of CARA traders, $\pi_{U,0}$ stands for their aggregate initial position and $\alpha_{U}$ denotes their aggregate risk tolerance. Also, one may allow the liquidity providers to have initial endowment $\pi_{N,0}\neq 0$, but this could just be absorbed into the supply $\Pi$. As such, we take $\pi_{N,0}=0$.
@@
By Pareto optimality of the initial endowments we mean that agents’ marginal utility from endowed wealth is proportional to the density
(3) $\frac{d{\mathbb{Q}}_{0}}{d\mathbb{P}}=\frac{e^{-\widehat{\Pi}^{\prime}X}}{\mathbb{E}\left[e^{-\widehat{\Pi}^{\prime}X}\right]}.$
We will use ${\mathbb{Q}}_{0}$ below when expressing equilibrium prices.
At time 00 when the signal arrives, $I$ and $U$, using their respective information sets, choose positions $\pi_{I},\pi_{U}$ to take in the risky asset, financing this choice by trading in the riskless asset. Writing the to-be-determined equilibrium price as $p$, the terminal wealth is
$\mathcal{W}^{\pi_{i}}:=\pi_{i,0}p+\pi_{i}(X-p);\qquad i\in\left\{I,U\right\}.$
The equilibrium clearing condition is $\Pi=\widehat{\pi}_{I}+\widehat{\pi}_{U}+Z_{N}$, where $\widehat{\pi}_{I},\widehat{\pi}_{U}$ are the optimal positions for $I$ and $U$, and $Z_{N}$ the noise trader demand. As it is more natural to present results for risk-aversion adjusted strategies, we write
$\psi_{i}:=\frac{\pi_{i}}{\alpha_{i}};\qquad i\in\left\{I,U\right\}.$
With this notation the clearing condition is
@@
(5) $\mathcal{W}^{\psi_{i}}:=\frac{1}{\alpha_{i}}\mathcal{W}^{\alpha_{i}\psi_{i}}=\widehat{\Pi}p+\psi_{i}(X-p);\qquad i\in\left\{I,U\right\}.$
Equilibrium Construction
We now construct the price impact equilibrium. At the end of the section we will also summarize the associated price-taking equilibrium (which is known in the literature, dating to Grossman and Stiglitz (1980)), and consider when there is no private information. These latter summaries are made with an eye towards the comparison results of Sections 2 and 3. Proofs of all results are in Appendix B (for the general market model).
Consider when the insider perceives, and hence looks to exploit, her market power, and internalizes her price impact through trading. As the insider’s private information is partially revealed to the market through her demand (which affects the uniformed trader’s optimal demand and hence the equilibrium price), by accounting for her ability to impact the market, she controls the signal revealed to the market. Therefore we expect the insider’s internalization of impact to affect not only equilibrium prices, but also equilibrium information transmission. By contrast, we assume the uniformed agent takes prices, and hence time 0 information, as given. We believe this setting is reasonable, as the uniformed trader is a representative agent for a mass of “small” uniformed traders who do not impact the market.
@@
As is common in the literature, we seek a linear impact equilibrium. In other words, the insider perceives that if she changes her position from $\pi_{I,0}=\alpha_{I}\psi_{I,0}$ to $\pi_{I}=\alpha_{I}\psi_{I}$, then the price will be an affine function of her trade combined with the noise trader demand,
(6) $p_{\iota}(\psi_{I},Z_{N})=V+M\left(\psi_{I}-\psi_{I,0}+\frac{Z_{N}}{\alpha_{I}}\right),$
for constants $V,M$ that are determined in equilibrium [^8] The affine structure could also be deduced by the insider if she first examined the equilibrium structure in the price taking case - see Remark 1.9 below. However, as affine impact is so common, we take it as a primitive., and where throughout we use the subscript “$\iota$” to stand for “impact”. Next, following the analysis of Rochet and Vila (1994) we assume the insider can see both her private signal and, for a given trade $\psi$, the price $p_{\iota}$ of (6). This will imply that the noise trader demand $Z_{N}$ is revealed to the insider through her signal and the time 00 price. [^9] This is in contrast to Spiegel and Subrahmanyam (1992) and Subrahmanyam (1991) and leads to a different equilibrium, but we believe it is a realistic assumption, given that the equilibrium price (as will be shown) is linear in the signal and noise.
To make this assumption precise, note that $G$ and $Z_{N}$ are the only random quantities revealed at time 00, and hence every insider strategy must be known using the information generated by $G$ and $Z_{N}$. Therefore, if the insider uses a strategy $\psi$ which reveals the noise trader demand $Z_{N}$ through the price, it must also be that $Z_{N}$ is known given the information generated by $G$ and $p_{\iota}(\psi,Z_{N})$. As such, we define set of acceptable trading strategies for the insider to be
Here, “$\psi\in\sigma(G,H)$” means that $\psi$ is $\sigma(G,H)$ measurable, and note that for any $\psi\in{\mathcal{A}}_{I}$, one has $\sigma(G,p_{\iota}(\psi,Z_{N}))=\sigma(G,Z_{N})$.
As the insider knows the noise demand, the public quantity $\psi+Z_{N}/\alpha_{I}$ is (partially) controlled by insider, and its effect on equilibrium can be understood as her price impact. Indeed, when the uniformed trader acts optimally as price taker, the price will take the affine form in (6). As the price is public, the combined insider and noise trader demand changes the public information set, effectively altering the precision of a public signal (taking the same form as in (1) just with a different noise term). Therefore, when the insider accounts for her price impact, the uniformed trader also recognizes that the precision of the public signal will change with the insider’s policy. This is in contrast to the price-taking case and in fact implies the public signal is always of a lower quality when the insider internalizes impact (as we prove in Proposition 2.1).
Identifying the equilibrium is equivalent to identifying $M$ and $V$ which clear the market. The first step is to characterize the insider’s optimal demand for any fixed $M$ and $V$, which is the solution to
@@
(8) $\begin{split}&\inf_{\psi\in{\mathcal{A}}_{I}}\mathbb{E}\left[e^{-\psi_{I,0}p_{\iota}\left(\psi,Z_{N}\right)-\psi\left(X-p_{\iota}\left(\psi,Z_{N}\right)\right)}\big{|}\sigma(G,Z_{N})\right].\end{split}$
To identify the optimal $\widehat{\psi}_{I}$, it is useful to express $M$ in terms of $1+p_{I}$, the precision of $X$ given the insider signal $G$
(9) $\begin{split}M=M(y)=\frac{y}{1+p_{I}}.\end{split}$
The dependence of $M$ on $y$ will force $V$ to depend on $y$ as well, so we write $V=V(y)$. As we will see, $01+2y>$ 0 ensures well-posedness of the minimization problem (8) but realistically we expect a positive impact function so that $y>0$. The following lemma identifies the insider’s optimal demand in terms of $y$.
Lemma 1.1.
Let $g,z\in\mathbb{R}$. On the set $\left\{G=g,Z_{N}=z\right\}$, the unique optimizer of (8) enforces
@@
(10) $\begin{split}\widehat{\psi}_{I,\iota}(g,z)-\psi_{I,0}+\frac{z}{\alpha_{I}}=\frac{p_{I}}{1+2y}\left(g+\Lambda_{\iota}(y)z\right)-\frac{\widehat{\Pi}+(1+p_{I})V(y)}{1+2y},\end{split}$
@@
(11) $\begin{split}\Lambda_{\iota}(y)=\frac{1+y}{\alpha_{I}p_{I}}.\end{split}$
Remark 1.2.
$\widehat{\psi}_{I,\iota}$ is in fact optimal among all functions $\psi$ (not just those in ${\mathcal{A}}_{I}$) because $\widehat{\psi}_{I,\iota}$ is invertible in $z$. Therefore, the class ${\mathcal{A}}_{I}$ poses no restriction.
Turning to the uninformed agent, the clearing condition (4) implies in equilibrium $\widehat{\psi}_{I,\iota}-\psi_{I,0}+Z_{N}/\alpha_{I}$ is publicly observable, and using (10) it is natural to define the market signal
(12) $H_{\iota}:=G+\Lambda_{\iota}Z_{N}=X+Z_{I}+\Lambda_{\iota}Z_{N},$
which is of the same form as the insider signal $G$, except with lower precision
(13) $p_{U,\iota}(y)=\left(\frac{1}{p_{I}}+\frac{p_{N}}{\Lambda^{2}_{\iota}(y)}\right)^{-1}=p_{I}\times\frac{\alpha_{I}^{2}p_{I}p_{N}}{(1+y)^{2}+\alpha_{I}^{2}p_{I}p_{N}}.$
By observing the price, the uniformed trader has time 00 information $\sigma(H_{\iota})$ and written as a function of $H_{\iota}$, the price is $p(H_{\iota})$ where
(14) $\begin{split}p_{\iota}(h_{\iota})&=M(y)\left(\widehat{\psi}_{I,\iota}-\psi_{I,0}+\frac{Z_{N}}{\alpha_{I}}\right)+V(y)=\frac{p_{I}y}{(1+p_{I})(1+2y)}\left(h_{\iota}-\frac{\widehat{\Pi}}{p_{I}}\right)+\frac{1+y}{1+2y}V(y).\end{split}$
As the uninformed is a price taker and the risk aversion adjusted initial wealth $\psi_{U,0}p_{\iota}(H_{\iota})$ factors out, his optimization problem is
$\begin{split}&\inf_{\psi\in\sigma(H_{\iota})}\mathbb{E}\left[e^{-\psi\left(X-p_{\iota}\left(H_{\iota}\right)\right)}\big{|}\sigma(H_{\iota})\right].\end{split}$
Similarly to Lemma 1.1 we obtain
Lemma 1.3.
The uninformed agent has risk-aversion adjusted optimal demand
(15) $\widehat{\psi}_{U,\iota}=p_{U,\iota}H_{\iota}-(1+p_{U,\iota})p_{\iota}(H_{\iota}).$
Relations (12)-(15) quantify the insider’s impact on the market signal $H_{\iota}$, the market signal precision $P_{U,\iota}$, the price $p_{\iota}(H_{\iota})$, and the uninformed trader’s optimal demand $\widehat{\psi}_{U,\iota}$. It is important to note that these impacts are not consistent with the price-taking equilibrium discussed below. When the insider internalizes her price impact, the uniformed trader takes it into account, and the coefficients of her affine demand function change (in fact, the reactions to the public signal and its precision are lower as we show in Section 2). This means that even when the insider submits her optimal price-impact demand in the price taking equilibrium, the market will not equilibrate to the price-impact price, as the uniformed trader’s optimal demand alters. We return to this point in Remark 1.10.
We now identify $y$ and $V(y)$ by enforcing the market-clearing condition (4) using the initial endowments of (2),
$\begin{split}(\alpha_{I}+\alpha_{U})\widehat{\Pi}&=\alpha_{I}\left(\widehat{\psi}_{I,\iota}(G,Z_{N})-\psi_{I,0}+\frac{Z_{N}}{\alpha_{I}}\right)+\alpha_{I}\widehat{\Pi}+\alpha_{U}\widehat{\psi}_{U,\iota}(H_{\iota}).\end{split}$
Taking into account (10), (14) and (15), this reduces to
(16) $\begin{split}(\alpha_{I}+\alpha_{U})\widehat{\Pi}&=\alpha_{I}\left(\frac{p_{I}}{1+2y}h_{\iota}-\frac{\widehat{\Pi}+(1+p_{I})V(y)}{1+2y}\right)+\alpha_{I}\widehat{\Pi}\\ &\qquad+\alpha_{U}\left(\left(p_{U,\iota}(y)-\frac{(1+p_{U,\iota})p_{I}y}{(1+p_{I})(1+2y)}\right)h_{\iota}+\frac{1+p_{U,\iota}}{1+2y}\left(\frac{y\widehat{\Pi}}{1+p_{I}}-V(y)\right)\right),\end{split}$
which will hold provided $y$ satisfies a certain cubic equation, as
Lemma 1.4.
Equation (16) holds provided $V(y)=-\widehat{\Pi}$ and $y$ solves
$\begin{split}0=\left(1+y\right)^{2}\left(1-\frac{\alpha_{U}y}{\alpha_{I}(1+p_{I})}\right)+\alpha_{I}^{2}p_{I}p_{N}\left(\frac{\alpha_{U}}{\alpha_{I}}y+\frac{\alpha_{I}+\alpha_{U}}{\alpha_{I}}\right).\end{split}$
As such, (positive) solutions to a cubic equation [^11] This differs from the quintic equation in Subrahmanyam (1991) when the insider does not see the noise trader demand. for $y$ are in one-to-one correspondence with equilibrium, with the positivity needed to ensure the pricing function is increasing in the combined trade $\widehat{\psi}_{I,\iota}+Z_{N}/\alpha_{I}$ (see (9)). Next, to ease the notation, define
@@
as the precision of $Z_{N}/\alpha_{I}$ and the insider’s proportion of the total risk tolerance respectively. The cubic equation becomes
and the following proposition shows
@@
There exists a unique solution $\widehat{y}>0$ to (18).
Given this proposition we establish equilibria in the main result of the section. To state it, recall the measure ${\mathbb{Q}}_{0}$ from (3) and define
(19) $p_{0}:=\mathbb{E}^{{\mathbb{Q}}_{0}}\left[X\right]=-\widehat{\Pi},$
as the equilibrium price, absent private information, given the initial endowments satisfy (2). By expressing quantities in terms of $p_{0}$ we provide an intuitive way to gauge how private information and price impact alter prices.
@@
$\widehat{y}>0$ from Proposition 1.5 induces a price-impact equilibrium. The market signal is $H_{\iota}$ from (12), and the equilibrium price is $p_{\iota}(H_{\iota})$ for the price function
(20) $\begin{split}p_{\iota}(h_{\iota})&=p_{0}+\frac{p_{I}\widehat{y}}{(1+p_{I})(1+2\widehat{y})}\left(h_{\iota}-p_{0}\right).\end{split}$
The insider has optimal policy $\widehat{\pi}_{I,\iota}=\widehat{\psi}_{I,\iota}(G,Z_{N})/\alpha_{I}$ where
$\widehat{\psi}_{I,\iota}(g,z)=\frac{1}{1+\widehat{y}}\left(p_{I}g-(1+p_{I})p_{\iota}(h_{\iota}(g,z))-\widehat{y}p_{0}\right).$
The uninformed agent has optimal policy $\widehat{\pi}_{U,\iota}=\widehat{\psi}_{U,\iota}(H_{\iota})/\alpha_{U}$ where from (15)
$\widehat{\psi}_{U,\iota}(h_{\iota})=p_{U,\iota}h_{\iota}-(1+p_{U,\iota})p_{\iota}(h_{\iota}).$
@@
For comparison purposes (i.e. to provide a benchmark case), herein we consider when all traders are price takers. As this result is well known (see Grossman and Stiglitz (1980)), we summarize the equilibrium structure in the following proposition. To state the proposition, assume there is a market signal $H$ revealed through the time 00 price $p=p(H)$, and both traders take $p(H)$ as given. The insider has time 0 information $\sigma(H,G)$ while the uninformed trader uses $\sigma(H)$. Using (5), the insider and uninformed trader’s optimal investment problems are respectively
@@
We say $(H,p(H))$ is a price-taking equilibrium if the clearing condition (4) holds for the optimal policies.
@@
(22) $H:=G+\Lambda Z_{N}=X+Z_{I}+\Lambda Z_{N},\qquad\Lambda=\frac{1}{\alpha_{I}p_{I}}.$
@@
(24) $\begin{split}p(h)&:=p_{0}+\frac{\alpha_{I}p_{I}+\alpha_{U}p_{U}}{\alpha_{I}(1+p_{I})+\alpha_{U}(1+p_{U})}(h-p_{0}).\end{split}$
The optimal policies for $I$ and $U$ are $\widehat{\psi}_{I}(G,Z_{N})$ and $\widehat{\psi}_{U}(H)$ respectively, where
(25) $\begin{split}\widehat{\psi}_{I}(g,z)&=p_{I}g-(p_{I}+1)p(h(g,z)),\qquad\widehat{\psi}_{U}(h)=p_{U}h-(p_{U}+1)p(h).\end{split}$
Remark 1.8.
Note that as $X$ is $\mathbb{P}$ independent of the noise terms $Z_{I},Z_{N}$ in (12), from the definition of ${\mathbb{Q}}_{0}$ in (3) $X$ is ${\mathbb{Q}}_{0}$ independent fo $Z_{I},Z_{N}$ as well, and $p_{0}=\mathbb{E}^{{\mathbb{Q}}_{0}}\left[X\right]=\mathbb{E}^{{\mathbb{Q}}_{0}}\left[H_{\iota}\right]=\mathbb{E}^{{\mathbb{Q}}_{0}}\left[p_{\iota}(H_{\iota})\right]$ and similarly $p_{0}=\mathbb{E}^{{\mathbb{Q}}_{0}}\left[p(H)\right]$. Therefore, on average under ${\mathbb{Q}}_{0}$, the prices with price impact/private information and without price impact/private information coincide.
Remark 1.9.
Our assumption of linear price impact is motivated by the price-taking case, where the impact is indeed linear. To see this, note from (22), (24) and (25), on the set $\left\{G=g,Z_{N}=z\right\}$, the combined trade of the insider and noise trader is
$\widehat{\psi}_{I}(g,z)-\psi_{I,0}+\frac{z}{\alpha_{I}}=p_{I}h(g,z)-(p_{I}+1)p(h(g,z))-\psi_{I,0}=\frac{\alpha_{U}(p_{I}-p_{U})}{\alpha_{I}p_{I}+\alpha_{U}p_{U}}\left(p(h)-p_{0}\right).$
This combined demand function, and the corresponding uniformed trader’s optimal demand, clears the market at the price $p(h)$. Solving this for $p(h)=p(h(g,z))$ we obtain
(26) $\begin{split}p(h(g,z))&=p_{0}+\frac{\alpha_{I}p_{I}+\alpha_{U}p_{U}}{\alpha_{U}(p_{I}-p_{U})}\left(\widehat{\psi}_{I}(g,z)-\psi_{I,0}+\frac{z}{\alpha_{I}}\right).\end{split}$
This is the reverse combined demand function at equilibrium, and indicates linear price impact. Indeed, even though the insider does not internalize impact in the price-taking case, in equilibrium it turns out the price is linearly impacted by her trade, combined with the noise trader’s demand. The price takes the form (6), where
(27) $V=p_{0},\qquad M=\frac{\alpha_{I}p_{I}+\alpha_{U}p_{U}}{\alpha_{U}(p_{I}-p_{U})}.$
Remark 1.10.
As discussed above, as long as the insider internalizes her price impact and the uniformed trader takes this into account, the price taking and price impact equilibria cannot coincide. In fact, even if the insider submits the demand which is optimal in the price-taking equilibrium, if she internalizes price impact, the market will not equilibrate to the price-taking equilibrium price. This would be the case if the uniformed trader did not perceive the change in market signal precision due to the insider’s demand (i.e., if he ignored the insider’s internalization of the price impact and assumed $p_{U,\iota}(y)=p_{U}$ for all $y$.)
On the other hand, there is a $y^{*}>0$ such that $M(y^{*})=y^{*}/(1+p_{I})$ coincides with $M$ from (27). For this $M(y^{*})$, if the insider used the price taking optimal demand $\widehat{\psi}_{I}$ from (25), it would reveal the same signal to the uniformed trader as in the price-taking equilibrium. This would lead to the same uniformed trader’s demand and hence the same clearing price as in price-taking equilibrium. However, when $M=M(y^{*})$ and $V=0$, $\widehat{\psi}_{I}$ from (25) is not optimal for the insider. This is because of (11), which identifies the optimal demand under any linear price impact is not consistent with this choice of $M$ and $V$.
No private signal (NS) equilibrium
The price-taking results of the previous section allow us to isolate the effects of price-impact in the presence of information asymmetry. In this section, we turn off the asymmetric information channel to analyze the effects due solely to internalization of price impact. We envision a situation where there is a market maker who is capable of moving prices, but who is not privately informed about the asset’s terminal payoff. She wants to move prices against a mass of (small) uninformed traders in a way to maximize her utility. Now, to formally establish equilibrium, one would have to repeat the analysis of this section removing the private signal, giving all agents the same information set. However, it turns out that the no-signal equilibrium (in both the price-impact and price-taking cases) coincides with the previously established equilibrium in the limit $p_{I}\to 0$. The resultant equilibrium prices and optimal positions are summarized in the following proposition, the proof of which is given in Appendix B.

### v2

@@ 1. The Equilibrium
We now present the model and construct the equilibrium. To both isolate the effects of price impact internalization and asymmetric information, and to keep the presentation/notation as simple as possible, we consider a simplified model with only one risky tradeable asset whose payoff has zero mean and unit variance, and where the initial endowments are at Pareto optimality absent private information [^6] Even though traders have CARA preferences, equilibrium quantities under price-impact depend on the initial endowments, as it is the insider’s trade off her initial position which impacts equilibrium prices. Assuming initial endowments are Pareto optimal absent private information turns off the channel where hedging demands based on the initial position affect equilibrium quantities, allowing us to isolate information and internalization effects. . The single asset model, and all the main results, are generalized in Appendix A to multiple dimensions with general means and covariances.
@@
The model has one period and all random quantities are defined on a probability space $(\Omega,\mathcal{F},\mathbb{P})$, where the beliefs $\mathbb{P}$ are common to all agents. The risky asset has terminal payoff $X\sim N(0,1)$ and positive supply $\Pi>0$. The risk-less asset is in zero net supply, with price normalized to $1$. Following the literature [^7] See Subrahmanyam (1991); Spiegel and Subrahmanyam (1992); Grossman (1976); Grossman and Stiglitz (1980) amongst many others., there is an insider $I$ who at time 00 obtains a private signal $G$ which is a noisy version of $X$, taking the form
$G=X+Z_{I};\qquad Z_{I}=\frac{1}{\sqrt{p_{I}}}\mathcal{E}_{I},$
where $\mathcal{E}_{I}\sim N(0,1)$ is independent of $X$ and $p_{I}>0$ is the signal precision. There is also a mass of uninformed traders who do not receive a private signal, but who in the equilibrium established below, will receive a market signal through the time 00 price. We assume the uninformed traders are price-takers, and following convention, we consider a representative agent $U$, hereafter called the uninformed trader. We assume both traders have exponential preferences with respective risk tolerances $\alpha_{I}$, $\alpha_{U}$. Lastly, there are liquidity providers (also called noise traders), denoted by $N$, with exogenous demand
@@
where $\mathcal{E}_{N}\sim N(0,1)$ is independent of both $X$ and $\mathcal{E}_{I}$. $p_{N}$ measures the noise trader demand precision, and, as mentioned in Kovalenkov and Vives (2014), can be thought of as a measure of the volume of the price inelastic demand. Traders $I$ and $U$ are endowed with (constant) share positions $\left\{\pi_{i,0}\right\}$ which are Pareto optimal absent private information [^8] Since the uninformed agent, being a price taker, can be seen as a representative agent for a group of CARA traders, $\pi_{U,0}$ stands for their aggregate initial position and $\alpha_{U}$ denotes their aggregate risk tolerance. Also, one may allow the liquidity providers to have initial endowment $\pi_{N,0}\neq 0$, but this could just be absorbed into the supply $\Pi$. As such, we take $\pi_{N,0}=0$.
@@
By Pareto optimality of the initial endowments we mean there is a measure ${\mathbb{Q}}_{0}$ equivalent to $\mathbb{P}$ such that agents’ marginal utility from endowed wealth is proportional to the density $d{\mathbb{Q}}_{0}/d\mathbb{P}$. Due to exponential preferences, the density takes the form
(2) $\frac{d{\mathbb{Q}}_{0}}{d\mathbb{P}}=\frac{e^{-\widehat{\Pi}X}}{\mathbb{E}\left[e^{-\widehat{\Pi}X}\right]}.$
At time 00 when the signal arrives, $I$ and $U$, using their respective information sets, choose positions $\pi_{I},\pi_{U}$ to take in the risky asset, financing this choice by trading in the riskless asset. As it is natural to present results for risk-aversion adjusted strategies (henceforth called a “strategy”), we write
$\psi_{i}:=\frac{\pi_{i}}{\alpha_{i}};\qquad i\in\left\{I,U\right\},$
so the clearing condition is
@@
(4) $\mathcal{W}^{\psi_{i}}:=\frac{1}{\alpha_{i}}\left(\pi_{i,0}p+\pi_{i}(X-p)\right)=\widehat{\Pi}p+\psi_{i}(X-p);\qquad i\in\left\{I,U\right\}.$
Price-impact equilibrium
@@
Due to the exponential-Gaussian structure, we expect a linear-impact equilibrium. In other words, the insider perceives that if she changes her position from $\pi_{I,0}=\alpha_{I}\psi_{I,0}$ to $\pi_{I}=\alpha_{I}\psi_{I}$, then the price will be an affine function of her trade combined with the noise trader demand,
(5) $p_{\iota}(\psi_{I},Z_{N})=V_{p,\iota}+M_{p,\iota}\left(\psi_{I}-\psi_{I,0}+\frac{Z_{N}}{\alpha_{I}}\right),$
[…]
@@
$\begin{split}&\inf_{\psi\in{\mathcal{A}}_{I}}\mathbb{E}\left[e^{-\psi_{I,0}p_{\iota}\left(\psi,Z_{N}\right)-\psi\left(X-p_{\iota}\left(\psi,Z_{N}\right)\right)}\big{|}\sigma(G,Z_{N})\right].\end{split}$
Due to linear impact and the exponential-Gaussian structure, the insider’s optimal policy $\widehat{\psi}_{I,\iota}$ is affine in the private signal $G$ and noise trader demand $Z_{N}$. Therefore, with an eye towards (5), we may write
$\widehat{\psi}_{I,\iota}(g,z)-\psi_{I,0}+\frac{z}{\alpha_{I}}=V_{I,\iota}+M_{I,\iota}\left(g+\Lambda_{\iota}z\right),$
for certain constants $V_{I,\iota},M_{I,\iota}$ and most importantly $\Lambda_{\iota}$ (explicitly given in (9) below), each of which depends upon the pricing coefficients $V_{p,\iota},M_{p,\iota}$. As the quantity on the left side above is publicly observable, this implies the market receives the signal
$H_{\iota}=G+\Lambda_{\iota}Z_{N},$
and from (5) we may view the price as $p_{\iota}(H_{\iota})$. Therefore, the uninformed agent’s acceptable policies ${\mathcal{A}}_{U}$ are functions $\psi=\psi(H_{\iota})$, and as the uninformed is a price taker (and also because the risk aversion adjusted initial wealth $\psi_{U,0}p_{\iota}(H_{\iota})$ factors out), his optimization problem is
@@
$\begin{split}&\inf_{\psi\in{\mathcal{A}}_{U}}\mathbb{E}\left[e^{-\psi\left(X-p_{\iota}\left(H_{\iota}\right)\right)}\big{|}\sigma(H_{\iota})\right].\end{split}$
Similarly to the insider, the uninformed agent’s optimal policy function is affine in the public signal $H_{\iota}$, with coefficients that also depend on $V_{p,\iota},M_{p,\iota}$. Therefore, if we enforce the market clearing condition (3), we end up with an equation of the form (which must hold with probability one)
$0=V_{H,\iota}(V_{p,\iota},M_{p,\iota})+M_{H,\iota}(V_{p,\iota},M_{p,\iota})\times H_{\iota},$
where we have made explicit the dependence of the coefficients on $V_{p,\iota},M_{p,\iota}$. Thus, provided one can find $V_{p,\iota},M_{p,\iota}$ such that $0=V_{H,\iota}=M_{H,\iota}$ there exists a linear price-impact equilibrium.
Similarly to Subrahmanyam (1991) where equilibrium quantities were governed by solutions to a fifth order equation, in our model equilibrium quantities are governed by positive solutions to the cubic equation
@@
are the precision of $Z_{N}/\alpha_{I}$ and the insider’s proportion of the total risk tolerance respectively. As a first step we have the following.
@@
There exists a unique strictly positive solution $\widehat{y}$ to (6).
Given this proposition, our main result establishes equilibrium. To state it, recall the measure ${\mathbb{Q}}_{0}$ from (2) and define
(8) $p_{{\mathbb{Q}}_{0}}:=\mathbb{E}^{{\mathbb{Q}}_{0}}\left[X\right]=-\widehat{\Pi},$
as the equilibrium price absent private information. By expressing quantities in terms of $p_{{\mathbb{Q}}_{0}}$ we provide an intuitive way to gauge how private information and price impact alter prices.
@@
$\widehat{y}$ from Proposition 1.1 induces a price-impact equilibrium. The market signal is
(9) $H_{\iota}=G+\Lambda_{\iota}Z_{N},\qquad\Lambda_{\iota}=\frac{1+\widehat{y}}{\alpha_{I}p_{I}}.$
$H_{\iota}$ is of the same form as the insider signal $G$, except with lower precision
(10) $p_{U,\iota}=\left(\frac{1}{p_{I}}+\frac{1}{p_{N}\Lambda^{2}_{\iota}}\right)^{-1}=p_{I}\times\frac{\kappa p_{I}}{(1+\widehat{y})^{2}+\kappa p_{I}}.$
The equilibrium price is $p_{\iota}(H_{\iota})$ for the price function
(11) $\begin{split}p_{\iota}(h_{\iota})&=p_{{\mathbb{Q}}_{0}}+\frac{p_{I}\widehat{y}}{(1+p_{I})(1+2\widehat{y})}\left(h_{\iota}-p_{{\mathbb{Q}}_{0}}\right).\end{split}$
Writing $H_{\iota}=h_{\iota}(G,Z_{N})$, the insider has optimal policy function
(12) $\begin{split}\widehat{\psi}_{I,\iota}(g,z)&=\frac{p_{I}}{1+\widehat{y}}g-\frac{1+p_{I}}{1+\widehat{y}}p_{\iota}(h_{\iota}(g,z))-\frac{\widehat{y}}{1+\widehat{y}}p_{{\mathbb{Q}}_{0}}.\end{split}$
The uninformed agent has optimal policy function
(13) $\widehat{\psi}_{U,\iota}(h_{\iota})=p_{U,\iota}h_{\iota}-(1+p_{U,\iota})p_{\iota}(h_{\iota}).$
Lastly, the coefficients in (5) are $M_{p,\iota}=\widehat{y}/(1+p_{I})$ and $V_{p,\iota}=-\widehat{\Pi}=p_{{\mathbb{Q}}_{0}}$.
@@
For comparison purposes, herein we consider the case where all agents are price takers. As this result is well known (see Grossman and Stiglitz (1980)), we summarize the equilibrium structure in the following proposition. To state it, assume there is a market signal $H$ revealed through the time 00 price $p=p(H)$, and both traders take $p(H)$ as given. The insider has time 0 information $\sigma(H,G)$ while the uninformed trader uses $\sigma(H)$. Using (4), the insider and uninformed trader’s optimal investment problems are respectively
@@
We say $(H,p(H))$ is a price-taking equilibrium if the clearing condition (3) holds for the optimal policies. In the proposition below, note the similarities to Theorem 1.2. This will form the basis of our comparison results in the next section.
@@
(15) $H:=G+\Lambda Z_{N},\qquad\Lambda=\frac{1}{\alpha_{I}p_{I}}.$
@@
(17) $\begin{split}p(h)&:=p_{{\mathbb{Q}}_{0}}+\frac{\alpha_{I}p_{I}+\alpha_{U}p_{U}}{\alpha_{I}(1+p_{I})+\alpha_{U}(1+p_{U})}(h-p_{{\mathbb{Q}}_{0}}).\end{split}$
Writing $H=h(G,Z_{N})$, the insider has optimal policy function
(18) $\begin{split}\widehat{\psi}_{I}(g,z)&=p_{I}g-(1+p_{I})p(h(g,z)).\end{split}$
The uninformed agent has optimal policy function
(19) $\widehat{\psi}_{U}(h)=p_{U}h-(1+p_{U})p(h).$
No private signal equilibria
Lastly, we provide results absent private information, where the only uncertainty at time 00 is the noise trader demand $Z_{N}$. We turn off the asymmetric information channel to analyze effects due solely to internalization of price impact. [^11] Qualitatively, the no signal limit corresponds to when there is a market maker who is capable of moving prices, but who is not privately informed about the asset’s terminal payoff. She aims to move prices against a mass of (small) uninformed traders in a way to maximize her utility. Here, it turns out that the no-signal equilibria (in both the price-impact and price-taking cases) coincide with the previously established equilibria in the limit $p_{I}\to 0$. The resultant prices and optimal positions are summarized in the following proposition, the proof of which is given in Appendix B.2.